# SeSE: A Structural Information-Guided Uncertainty Quantification Framework for Hallucination Detection in LLMs

### v2

@@ VIII-B Hierarchical Abstracting
Figure 10 illustrates the construction of a $2$-dimensional hierarchical abstraction. Specifically, consider a semantic space represented by an irreducible non-negative matrix $A_{n\times n}$. Let $V=\{1,2,\cdots,n\}$, and let $\mathcal{T}$ be an encoding tree for $A$. Assume $\alpha$ and $\beta$ are two leaf nodes in $\mathcal{T}$ that share a common parent node $\gamma$, i.e., $\alpha^{-}=\beta^{-}=\gamma$. The steps for adjusting the encoding tree using the merging operator of leaf nodes $\alpha$ and $\beta$ are as follows [31]:
@@
Assume $\alpha$ and $\beta$ are two arbitrary nodes in $\mathcal{T}$ that share a common parent node $\gamma$, i.e., $\alpha^{-}=\beta^{-}=\gamma$. The steps for adjusting the encoding tree using the combining operator of $\alpha$ and $\beta$ are as follows:
@@
The 2-dimensional encoding tree could then be optimized to the needed $K$-dimension by continuing to greedily and iteratively apply merging and combining operators.
Figure 10: Illustration of the hierarchical abstracting construction.
VIII-C Computational Complexity and Cost
This section analyzes the time complexity and computational cost of our SeSE framework to evaluate its feasibility.
VIII-C1 Computational Cost
In terms of computational cost, model inference is involved only in the response sampling and semantic graph construction stages. The first step requires obtaining $N$ samples, thus the original LLM needs to perform $N$ inference operations. The second step requires using an NLI model to perform $N^{2}$ bidirectional inferences to build a directed graph. Since the parameter scale of NLI models is typically much smaller than that of LLMs (e.g., the DeBERTa-large-MNLI model used in this paper has only 1.5 billion parameters), the overall computational cost is primarily determined by the number of LLM inferences in the first step. Our hyperparameter sensitivity experiments show that a relatively small sample size $N=5$ is sufficient to achieve strong performance, thus the actual computational cost is low. AAlthough the sampling process introduces a minor additional expense, this cost is negligible compared with supervised methods requiring model retraining, white-box approaches dependent on internal model states, or techniques that rely on large and costly knowledge bases. By relying solely on sampled responses under black-box conditions, SeSE is plug-and-play and can adapt to both open-source and closed-source models, offering broader applicability.
VIII-C2 Time Complexity
@@
The overall time complexity of the hierarchical abstraction is $O(n^{2}+n+m\cdot\log_{2}n)$, where $n=N=|V|$ denotes the number of vertices in the semantic graph, and $m=|E|\approx\dfrac{kn}{2}$ (where $k$ represents the out-degree per vertex in the kkNN graph) indicates the number of edges retained after sparsification. The graph construction phase exhibits a time complexity of $O(n^{2}+n)$: constructing the complete graph requires $O(n^{2})$, and the sparsification operation requires $O(n)$. According to the analysis of [28], the optimization process of the high-dimensional encoding tree via stretch and compression operators contributes a time complexity of $O(m\cdot\log_{2}n)$ .
VIII-D Additional Experimental Details and Analysis
In this section, we provide additional experimental results.
VIII-D1 Hardware and Resources
@@
In terms of computing resources, as it is necessary to sample generations from LLMs to model the semantic space, our experiments require one or more GPUs to accelerate LLMs inference. Without GPU support, reproducing the results within a reasonable timeframe is infeasible. For sentence-length generation tasks, we use the GPT‑5‑mini model accessed through the OpenAI API to evaluate accuracy. As OpenAI’s pricing is based on the number of input and output tokens, the cost of reproducing our experiments varies with configuration, typically averaging around 1\$-5\$ per run. The concurrent experiments are conducted on two NVIDIA RTX PRO 6000 graphics cards, each with 96 GB of memory. Depending on the model and experimental setup, the generation phase for each scenario requires between 2 and 24 hours. Our experimental procedure involves first generating responses for all dataset-model pairs, followed by the computation of various uncertainty metrics. Model outputs are not regenerated across runs; instead, only the corresponding uncertainty metrics are recalculated.
VIII-D2 Assessing accuracy of automated ground-truth evaluations
@@
The F1 score is the harmonic mean of precision and recall of the lexical overlap between the reference answer and the generated answer. It is widely used to evaluate fixed-answer generation tasks. However, this metric exhibits obvious limitations in sentence-length, free-form text generation: lexical overlap between LLMs responses and the short reference answers may be unreasonably low, rendering the F1 score ineffective. Therefore, our study leverages the natural language understanding capabilities of LLMs rather than relying on simple lexical matching. We employ GPT-5-mini to assess semantic equivalence between LLM-generated answers and reference answers.
@@
To verify the reliability of our automated factual evaluation, we manually inspect 500 questions (100 from each of five experimental datasets) and analyze the sentence-length answers generated by the models. We focus on the concordance between human and automated methods, rather than the correctness of the evaluations. Table IV presents consistency statistics between automated evaluation methods and human reviewer judgments. The table shows that the agreement rate between the two human assessors (95%) closely approximates their average agreement rate with GPT-5-mini (94%). Although GPT-4o and Qwen-3-32B perform slightly worse, we select GPT-5-mini for the results presented in this paper, as it provides the best factual estimates.
TABLE IV: Assessing automated ground-truth evaluators.
@@
VIII-D3 Assessing entailment estimator in long-form generation
The bipartite graph construction involves entailment judgments between long-form responses and short claims. Figure 11 shows the ablation results of different entailment estimators. Conventional BERTScore and DeBERTa perform poorly, and GPT-5 does not offer notable advantages compared to GPT-5-mini. Consequently, we choose GPT-5-mini for its comparable performance and greater cost-effectiveness. The experiment is conducted on the PopQA using Gemini-2.5-Flash.
@@ VIII-D3 Assessing entailment estimator in long-form generation
VIII-D4 Detailed results of hyperparameter sensitivity
Figure 12: Statistics of semantic cluster numbers across datasets used in the sentence-length experiment. All plots are based on generations of Llama-3.1-70B.
Figure 13: AUROC performance across different sample sizes in sentence-length experiments.
Figure 14: AURAC performance across different sample sizes in long-form experiments.
In Figure 12, we present the empirical statistics of semantic cluster counts across datasets used in the sentence-length experiment. As illustrated, the semantic spaces of simpler datasets such as TriviaQA and SVAMP usually contain only one or two clusters, reflecting the fact that large language models generally answer these questions correctly. In contrast, more challenging datasets like SQuAD exhibit much higher semantic complexity: approximately 30% of the questions correspond to semantic spaces with six or more clusters, indicating more intricate and disordered semantic structures. From an information-theoretic perspective, such complex semantic spaces are difficult to compress. Accurately describing them therefore requires more information (bits), i.e., constructing deeper encoding trees (larger $K$) to effectively quantify their inherent uncertainty.
In Figures 1314, we report the unaggregated hyperparameter sensitivity results of the sampling size $N$ in sentence-length experiments. Figure 15 shows the unaggregated hyperparameter sensitivity results in the long-form experiments.
@@ VIII-D4 Detailed results of hyperparameter sensitivity
VIII-E Details of Case Study
Detailed information about the examples used in the case study is presented in Table V and Figure 16, based on the generations of Llama-3.1-70B on the SQuAD dataset. As shown in Figure 16, the semantic spaces of the two cases appear superficially similar: both can be divided into five semantic clusters, and the Shannon entropy values of these clusters are nearly identical. Since SE can only consider the probability distribution of semantic equivalence relations, it underestimates the uncertainty discrepancy between them, potentially leading to misjudgment. In contrast, SeSE provides a more precise assessment of uncertainty by incorporating more expressive semantic structural information. This case illustrates a representative scenario in our analysis: although the semantic spaces of the answers appear similar, intrinsic structural differences persist, and SeSE demonstrates superior performance to SE under such conditions.
TABLE V: Examples from SQuAD that are generated by Llama-3.1-70B.
Hallucination case
Question: What spacecraft contained data to determine the oxygen content of the Sun?
Answer: Genesis spacecraft
Sampled Responses:
1. Mariner 2 contained data to determine the oxygen content of the Sun.
2. Helios 1 and 2 contained data to determine the oxygen content of the Sun.
3. The Helios spacecraft contained data to determine the oxygen content of the Sun.
4. The Mars Global Surveyor contained data to determine the oxygen content of the Sun.
5. The Helios spacecraft contained data to determine the oxygen content of the Sun.
6. The Apollo 15 command module contained data to determine the oxygen content of the Sun.
7. The ACRIM-1 experiment was used to determine the oxygen content of the Sun.
8. Helios 1 contained data to determine the oxygen content of the Sun.
9. Helios contained data to determine the oxygen content of the Sun.
10. The Helios spacecraft contained data to determine the oxygen content of the Sun.
Non-hallucination case
Question: What gearing was used on steam turbine marine engines in the 20th century?
Answer:reduction
Sampled Responses:
1. In the 20th century, steam turbine marine engines often used reduction gearing to connect the turbine to the propeller shaft.
2. In the 20th century, single or double helical gearing was used on steam turbine marine engines.
3. In the 20th century, the gearing on steam turbine marine engines typically had pinions with 8 to 13 teeth and wheels with 60 to 100 teeth, with gear
ratios usually around 5:1 to 7:1.
4. In the 20th century, most steam turbine marine engines used single reduction gearing, although a few used double reduction or triple reduction gears.
5. In the 20th century, steam turbine marine engines typically used one or more sets of reduction gears, along with one or more steam turbines, including
power turbines and condensing turbines, to drive the propeller shaft.
6. In the 20th century, steam turbine marine engines used reduction gearing and epicyclic gearing.
7. In the 20th century, most steam turbine marine engines used single-reduction gearing, although some triple-expansion and other reciprocating steam
engines used double-reduction gearing.
8. In the 20th century, most steam turbine marine engines used single reduction gearing with double helical gears on a single pinion; some US Navy ships
used triple reduction gearing, some merchant ships and submarines used a combination of gearing and a fluid coupling, the British Royal Navy used double
reduction gearing, and some high-speed, high-powered ships used triple reduction gearing or single reduction with a fluid coupling, while some smaller ships
and submarines used single reduction without a fluid coupling, making single reduction with double helical gears the most common gearing.
9. In the 20th century, steam turbine marine engines used epicyclic gearing.
10. In the 20th century, gearing on steam turbine marine engines typically involved a combination of a high-pressure turbine and a low-pressure turbine
connected to a reduction gear that drove the propeller shaft.
(a) Heat map of the adjacency matrix that models the semantic space of the used hallucination cases.
(b) Heat map of the adjacency matrix that models the semantic space of the used non-hallucination cases.
Figure 16: The visualization of semantic spaces that is used for case study.
VIII-F Prompt Details
Here we mark placeholders with the orange color.
VIII-F1 Response sampling prompt
For sentence-length experiments, we use the following generation template:
  • Answer the following question concisely in one complete sentence.

### v3

@@ Appendix C Details of Hierarchical Abstraction
Figure 6 illustrates the construction of a $2$-dimensional hierarchical abstraction. Specifically, consider a semantic space represented by a strongly connected graph $G=(V,W,E)$ with non-negative normalized edge weights, which is equally to an irreducible non-negative matrix $A_{n\times n}$. Let $V=\{1,2,\cdots,n\}$, and let $\mathcal{T}$ be an encoding tree for $A$. Assume $\alpha$ and $\beta$ are two leaf nodes in $\mathcal{T}$ that share a common parent node $\gamma$, i.e., $\alpha^{-}=\beta^{-}=\gamma$. The steps for optimizing the encoding tree using the merging operator of leaf nodes $\alpha$ and $\beta$ are as follows:
@@
We define $g_{\alpha,\beta}$ as the total weight of edges connecting vertices in $V_{\alpha}$ to vertices in $V_{\beta}$ as follows:
$g_{\alpha,\beta}=\sum_{v_{i}\in V_{\alpha}}\sum_{v_{j}\in V_{\beta}}\pi(v_{i})\cdot W(v_{i},v_{j}).$ (18)
The entropy variation caused by a single merge operation on sibling nodes $\alpha,\beta\in T$ is denoted as $\Delta^{op_{mer}}_{\alpha,\beta}$ and is calculated as follows:
$\begin{aligned} \Delta^{op_{mer}}_{\alpha,\beta}&=\left[H^{\mathcal{T}}\left(G;\alpha\right)+H^{\mathcal{T}}\left(G;\beta\right)+\sum_{i}^{L_{\alpha}}H^{\mathcal{T}}\left(G;\alpha_{i}\right)+\sum_{i}^{L_{\beta}}H^{\mathcal{T}}\left(G;\beta_{i}\right)\right]-\left[H^{\mathcal{T}^{\prime}}\left(G;\mu_{mer}\right)+\sum_{i}^{L_{\alpha}}H^{\mathcal{T}^{\prime}}\left(G;\alpha_{i}\right)+\sum_{i}^{L_{\beta}}H^{\mathcal{T}^{\prime}}\left(G;\beta_{i}\right)\right]\\ &=\frac{g_{\alpha}-\sum_{i}^{L_{\alpha}}g_{\alpha_{i}}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\mu_{mer}}}{\mathcal{V}_{\alpha}}+\frac{g_{\beta}-\sum_{i}^{L_{\beta}}g_{\beta_{i}}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\mu_{mer}}}{\mathcal{V}_{\beta}}+\frac{g_{\alpha,\beta}+g_{\beta,\alpha}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\alpha^{-}}}{\mathcal{V}_{\mu_{mer}}}\\ &=\frac{g_{\alpha,\beta}+g_{\beta,\alpha}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\alpha^{-}}}{\mathcal{V}_{\mu_{mer}}}-\frac{\sum_{i\neq j}^{L_{\alpha}}g_{\alpha_{i},\alpha_{j}}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\mu_{mer}}}{\mathcal{V}_{\alpha}}-\frac{\sum_{i\neq j}^{L_{\beta}}g_{\beta_{i},\beta_{j}}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\mu_{mer}}}{\mathcal{V}_{\beta}},\end{aligned}$
where $\mathcal{T}^{\prime}$ represents the tree after the merge operation, $L_{\alpha}$ denotes the number of child nodes of $\alpha$, and $\mu_{mer}$ is the newly added node created by the merge operation.
Assume $\alpha$ and $\beta$ are two arbitrary nodes in $\mathcal{T}$ that share a common parent node $\gamma$, i.e., $\alpha^{-}=\beta^{-}=\gamma$. The steps for optimizing the encoding tree using the combining operator of $\alpha$ and $\beta$ are as follows:
@@
The entropy variation caused by a single combine on sibling nodes $\alpha,\beta\in T$ is denoted as $\Delta^{op_{com}}_{\alpha,\beta}$ and is calculated as follows:
$\Delta^{op_{com}}_{\alpha,\beta}$ $=\left[H^{\mathcal{T}}\left(G;\alpha\right)+H^{\mathcal{T}}\left(G;\beta\right)\right]-\left[H^{\mathcal{T}^{\prime}}\left(G;\mu_{com}\right)+H^{\mathcal{T}^{\prime}}\left(G;\alpha\right)+H^{\mathcal{T}^{\prime}}\left(G;\beta\right)\right]$ (20)
$=\frac{g_{\alpha}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\alpha^{-}}}{\mathcal{V}_{\mu_{com}}}+\frac{g_{\beta}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\alpha^{-}}}{\mathcal{V}_{\mu_{com}}}-\frac{g_{\mu_{com}}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\alpha^{-}}}{\mathcal{V}_{\mu_{com}}}$
$=\frac{g_{\alpha}+g_{\beta}-g_{\mu_{com}}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\alpha^{-}}}{\mathcal{V}_{\mu_{com}}}$
$=\frac{g_{\alpha,\beta}+g_{\beta,\alpha}}{\operatorname{vol}\left(G\right)}\cdot\log_{2}\frac{\mathcal{V}_{\alpha^{-}}}{\mathcal{V}_{\mu_{com}}},$
@@
where $\mu_{com}$ is the newly added node via the combine operation. The $2$-dimensional encoding tree could then be optimized to the needed $K$-dimension by continuing to greedily and iteratively apply merging and combining operators.
Time Complexity of Hierarchical Abstraction Construction
@@
The overall time complexity of the hierarchical abstraction construction is $O(n^{2}+m\cdot\log_{2}n)$ (Step 2 and Step 3 in subsection 3.1), where $n=|V|$ denotes the number of vertices in the semantic graph, and $m=|E|$ indicates the number of edges. The graph construction phase exhibits a time complexity of $O(n^{2})$. According to the analysis of Pan et al. [2025], the optimization process of the high-dimensional encoding tree via merging and combining operators contributes a time complexity of $O(m\cdot\log_{2}n)$.
Appendix D Additional Experimental Details and Analysis
D.1 Hardware and Resources
@@
In terms of computing resources, as it is necessary to sample generations from LLMs to model the semantic space, our experiments require one or more GPUs to accelerate LLMs inference. Without GPU support, reproducing the results within a reasonable timeframe is infeasible. For short-form generation tasks, we use the GPT-5-mini model accessed through the OpenAI API to evaluate accuracy. As OpenAI’s pricing is based on the number of input and output tokens, the cost of reproducing our experiments varies with configuration, typically averaging around 1\$-5\$ per run. The concurrent experiments are conducted on two NVIDIA RTX PRO 6000 graphics cards, each with 96 GB of memory. Depending on the model size and experimental setup, the generation phase for each scenario requires between 2 and 24 hours. Our experimental procedure involves first generating responses for all dataset-model pairs, followed by the computation of various uncertainty metrics. Model outputs are not regenerated across runs; instead, only the corresponding uncertainty metrics are recalculated.
D.2 Comparison of Computational Resource Consumption
In this subsection, we analyze the computational resource consumption of various UQ methods. Given the vast parameter size of LLMs, their inference costs are significantly more intensive than other components; therefore, our analysis primarily focuses on LLM inference consumption. Besides the white-box method Length-normalized Predictive Entropy (LN-PE), the few-shot prompting method P(True), and the supervised training method Embedding Regression (ER), all other baselines require sampling $N$ possible answers, incurring the same consumption. To ensure accuracy in semantic clustering for SE and self-consistency assessment for SC, following the original implementations, we employ GPT-5-mini for entailment prediction. In the worst-case scenario, this necessitates $N^{2}$ LLM inference calls. Graph-based methods (KLE and SGD) utilize lightweight NLI models to compute entailment scores between responses, which is identical to SeSE. In summary, SeSE significantly reduces computational costs compared to SE and SC while achieving superior performance. When compared to graph-based methods, SeSE delivers better results while maintaining equivalent resource consumption. Although the sampling process increases the generation cost, SeSE avoids the limitations of supervised methods (ER) that require retraining for new models and tasks, or white-box methods (LN-PE, P(True)) that depend on LLM internal states. Moreover, in safety-critical tasks, the potential cost of a hallucination should outweigh the cost of sampling multiple answers; therefore, reliable uncertainty quantification through SeSE should always be worthwhile.
D.3 Generalization Analysis
Figure 7(a) illustrates the hallucination rates of used LLMs on each dataset. As can be seen, our experiments cover a range of LLMs with varying hallucination levels across diverse generation tasks. The values plotted in Figure 7(b) represent the aggregate AUROC scores over five LLMs. Embedding regression is a representative supervised approach that utilizing a trained logistic regression classifier to predict answer correctness. P(True) serves as an “in-context" supervised method that adapts to specific tasks through few-shot demonstrations in the prompt. As indicated by the light red and light purple bars, both P(True) and ER suffer substantial performance degradation when the data distribution shifts between training and testing. In contrast, as an “off-the-shelf” method, SeSE consistently outperforms both supervised and entropy-based baselines on in-distribution and out-of-distribution (OOD) datasets, demonstrating significant generalization and practical utility potential. Such generalization is especially important for UQ, as real-world scenarios frequently involve distribution shifts between training and deployment phases, and reliable UQ methods should perform well across different scenarios.
(a) Hallucination rate of LLMs across used datasets.
(b) OOD performance comparison among entropy-based methods.
Figure 7: Hallucination rate of used LLMs in different domains and performance comparison of entropy-based methods in OOD datasets. OOD represents the method is evaluated on out-of-distribution datasets.
D.4 Comparison with Graph Centrality Metrics
Regarding granular uncertainty estimation in long-form generation, the most closely related work to ours is Graph Uncertainty Jiang et al. [2024], which utilizes the negative centrality scores of claim nodes within a claim-response bipartite graph as the uncertainty metric. However, the centrality metrics employed by [Jiang et al., 2024] fail to capture the richer semantic topological structures Li and Pan [2016] and offer limited interpretability. Here, we compare SeSE with several widely used graph centrality measures, including betweenness, eigenvector, PageRank, and closeness Lin et al. [2024]. The specific definitions of these metrics are detailed in Table 3. We follow the same setup in main experiments. As illustrated in Figure 8, SeSE consistently outperforms these graph centrality metrics, demonstrating its superior ability to identify central nodes within the claim-response graph.
Figure 8: SeSE significantly outperforms benchmark graph centrality metrics in Table 3.
Table 3: Graph centrality metrics with their formulas and explanations.
Metric Formula Brief Explanation
Eigenvector $C_{Eigv}(v)=\frac{1}{\lambda}\sum_{u\in N(v)}A_{vu}C_{Eigv}(u)$ Evaluates the influence of node $v$ based on the importance of its neighbors $N(v)$. $A_{vu}$ is adjacency matrix entry. $\lambda$ is the largest eigenvalue of $A$.
D.5 Assessing accuracy of automated ground-truth evaluations
@@
The F1 score is the harmonic mean of precision and recall of the lexical overlap between the reference answer and the generated answer. It is widely used to evaluate fixed-answer generation tasks. However, this metric exhibits obvious limitations in short-form, free-form text generation: lexical overlap between LLMs responses and the short reference answers may be unreasonably low, rendering the F1 score ineffective. Therefore, our study leverages the natural language understanding capabilities of LLMs rather than relying on simple lexical matching. We employ GPT-5-mini to assess semantic equivalence between LLM-generated answers and reference answers.
@@
To verify the reliability of our automated factual evaluation, we manually inspect 500 questions (100 from each of five experimental datasets) and analyze the short-form answers generated by the models. We focus on the concordance between human and automated methods, rather than the correctness of the evaluations. Table 4 presents consistency statistics between automated evaluation methods and human reviewer judgments. The table shows that the agreement rate between the two human assessors (95%) closely approximates their average agreement rate with GPT-5-mini (94%). Although GPT-4o and Qwen-3-32B perform slightly worse, we select GPT-5-mini for the results presented in this paper, as it provides the best factual estimates.
Table 4: Assessing automated ground-truth evaluators.
@@
D.6 Assessing entailment estimator in long-form generation
The bipartite graph construction involves entailment judgments between long-form responses and short claims. Figure 9 shows the ablation results of different entailment estimators. Conventional BERTScore and DeBERTa perform poorly, and GPT-5 does not offer notable advantages compared to GPT-5-mini. Consequently, we choose GPT-5-mini for its comparable performance and greater cost-effectiveness. The experiment is conducted on the PopQA using Gemini-3-Flash.
@@ D.6 Assessing entailment estimator in long-form generation
Figure 11: Statistics of semantic cluster numbers across datasets used in the short-form experiment. All plots are based on generations of Llama-3.1-70B.
Figure 12: AUROC performance across different sample sizes in short-form experiments.
Figure 13: AURAC performance across different sample sizes in short-form experiments.
D.7 Detailed results of hyperparameter sensitivity
In Figure 11, we present the empirical statistics of semantic cluster counts across datasets used in the short-form experiment. As illustrated, the semantic spaces of simpler datasets such as TriviaQA and SVAMP usually contain only one or two clusters, reflecting the fact that large language models generally answer these questions correctly. In contrast, more challenging datasets like SQuAD exhibit much higher semantic complexity: approximately 30% of the questions correspond to semantic spaces with six or more clusters, indicating more intricate and disordered semantic structures. From an information-theoretic perspective, such complex semantic spaces are difficult to compress. Accurately describing them therefore requires more information (bits), i.e., constructing deeper encoding trees (larger $K$) to effectively quantify their inherent uncertainty. Figure 10 shows the unaggregated hyperparameter sensitivity results in the long-form experiments. In Figures 1213, we report the unaggregated hyperparameter sensitivity results of the sampling size $N$ in short-form experiments.
Appendix E Prompt Details
Here we mark placeholders with the blue color.
E.1 Response sampling prompt
We follow Farquhar et al. [2024] and use the following prompt template to obtain answers in short-form experiments, including both the greedily-decoded answer (to evaluate the model’s accuracy) and the randomly sampled responses (to measure the model’s uncertainty).
  • Answer the following question concisely in one sentence.